A generator is designed to produce a maximum emf of 180 Vwhile rotating with an angular speed of 3300 rpm . Each coilof the generator has an area of1.9×10−2 m2.
If the magnetic field used in the generatorhas a magnitude of 4.5×10−2 T , how many turns of wire are needed?
Answer #1

Maximum emf E = 180 V

Angular speed w = 3300 rpm

                           = 3300 * 2(pi) rad / 60 s

                           = 345.57 rad / s

Area   A= 1.9 * 10^-2 m ^ 2

Magnetic field B = 4.5 * 10 ^ -2 T

We know E = BANw

From this No.of truns N = E / ( BA w )

                                        = 609.2 ~ 610